In 1993,10,631,000 vehicles were registered in New Mexico, Maryland, and Georgia. New Mexico had 3,000 less than 2/5ths of the number in Maryland. Georgia has 52,000 less than 4 times the number in New Mexico. how many registered motor vehicles were there in each of these three states??

so far I have the formulas:

NM=(MD)(2/5) - 3,000
GA= (NM)(4) - 52,000
MD= 10,613,000 - (NM + GA)

Answer by rapaljer(4671)

It looks right so far.
Substitute the first formula for NM into the second equation:
GA = 4(NM) - 52000
GA = 4(2/5MD -3000) - 52000
GA = 8/5 MD - 12000 - 52000
GA = 8/5 MD - 64000

Now, substitute THIS GA and the NM into the third equation:
MD = 10613000 - (NM + GA)
MD = 10613000 - ( 2/5MD -3000 + 8/5 MD -64000)
MD = 10613000 - (10/5 MD - 67000)
MD = 10613000 - 2MD + 67000
MD = 10680000 - 2MD

Add +2MD to each side of the equation:
3MD = 10680000 Divide by 3
MD = 3560000
It's late! I have no idea if this is right or not! It's the UGLIEST problem I've seen in a long time! At least it came out even, right?? Ask me in the morning!! Oops! It IS morning!!

R^2 at SCC